6) A Superbowl Party has 8 Green Bay Packer fans, 10 New England Patriot fans, 6 Oakland Raiders fans, 9 Dallas Cowboy Fans, and 5 Pittsburg Steeler fans. If 4 fans are selected to go to the store to pick up more chips and dips, then compute the probability that at least one of them will be a Dallas Cowboy fan. 7) A school cdlub of 40 students must select a President, a Vice-President, a secretary, and a treasurer. How many ways can this be done, if no club member can hold more than one office. 40 P4 A21433100 40-4)L
Answer #1

6)

Total number of fans:

8+10+6+9+5 = 38

Number of ways of selecting 4 fans out of 38 is

38! c(38,4) 438-4 73815

There are 38 -9 = 29 fans other than Dallas Cowboy fans so number of ways of selecting 4 fans out of 29 is

291 4029-4)! C(29,4) = _ 23751

Number of ways of selecting at least one Dallas fan is

73815 - 23751 = 50064

Hence, the required probability is

p = 50064 / 73815 = 0.6782